OpenStudy (anonymous):

identify zeros of the function(x-8)(x-15)

OpenStudy (anonymous):

Using the zero product property \((x - 8)(x - 15) = 0\) if and only if either one of the factors = 0.

OpenStudy (anonymous):

Basically, if the product is going to be zero, either x - 8 = 0 or x - 15 = 0.

OpenStudy (anonymous):

Can you solve from there?